OpenStudy (anonymous):

simplify 3+1/4 over 5/6+2 to the 2nd power

OpenStudy (anonymous):

12 /4 +1/4 = 13/ 4 , 5/6+12/6 = 17/6 ===> 3+1/4 over 5/6+2 = 13/4 over 17/6 = 39/34